Wendy has $5,000.00 invested in a bank that pays 6.25% annually. How long will it take for her funds to triple?

Round your answer to two decimal places. For example, if your answer is $345.667 round as 345.67 and if your answer is .05718 or 5.718% round as 5.72.

Group of answer choices

16.49 years

22.11 years

18.12 years

13.95 years

16.67 years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

(3*5000)=5000*(1.0625)^n

3=(1.0625)^n

Taking log on both sides;

log 3=n*log (1.0625)

n=log 3/log (1.0625)

=18.12 years(Approx).